# Can Robots “Taste” Grapes? Estimating SSC with Simple RGB Sensors

### v1

@@ Acknowledgments
This work is part of a project that has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No.
101016906 – Project CANOPIES
@@
Appendix A Detailed dataset description
Data collection was part of a larger effort within the Canopies project, exploring different sensors across seasons. The first data collection was done in September-October 2021, using a smartphone camera and a Reflex camera. The second data collection took place in September 2022, utilizing a smartphone camera and a Realsense d435i. Although the d435i is an RGB-D device, we used it as a standard camera for °Brix and color estimation.
Table 5 summarizes the different types of data collected during the two seasons. The naming of the splits in the following tables follows a descriptive naming used during the development of the dataset and the code. For ease of reference, we renamed the splits with simpler tags ($T1$, $T2$, $T3$) to avoid bloated experiment descriptions. The match between the splits names given in Table 1 and 7 are the following:
[…]
The data organization allows for testing covariate shifts due to both sensor type and season. These atomic splits represent minimal groups sharing the same sensor, season, image type, and label characteristics. Using these atomic splits, we designed the experimental splits, summarized in Table 6 and Table 7.
Images: Smartphone images were acquired at a resolution of 3000x4000 pixels, in both JPG and raw formats, with a few exceptions. Reflex images, captured in raw format, have a resolution of 3900x2613 pixels. The d435i RGB images have a lower resolution of 1280x720 pixels. Raw images were corrected for white balance using a gray card (Fig.2). Not all images included a gray card due to collection errors, but these images were still used for comparison purposes. The full images were labeled with bounding boxes around the grape bunches and then cropped. The resulting cropped images have very different sizes since some images were close-ups while others contained multiple targets. The combination of season, sensor used, gray card presence, and other factors creates differences summarized in Table 5. The rows in the table represent atomic splits, i.e., splits with the same distribution. In contrast, between the splits, there is some form of covariate shift that we aim to study.
Datset Id Year Smartphone Reflex d435i Num images Num samples Expert GT Color GT
Type Jpg Raw Card Raw Card RGB IR Card
A 2021 motog8 yes no no no - no no - 100 100 no no
B 2021 motog8 no yes yes yes yes no no - 38 38 possible yes
C 2021 motog8 yes yes yes yes yes no no - 62 62 possible yes
D 2021 motog8 yes yes yes yes yes no no - 50 50 possible yes
E 2022 realmi yes yes no no - yes no no 1 2 no no
F 2022 realmi yes yes yes no - yes no yes 6 22 possible yes
G 2022 realmi yes yes no no - yes no yes 1 2 no no
H 2022 realmi yes yes yes no - yes no yes 4 19 possible yes
I 2022 motog8 yes yes yes no - yes yes yes 8 51 possible yes
Name Components Description
2021PhoneJPG A+C+D Contains all the jpg images collected with a smartphone in 2021, without white balance from the spits A, C, and D.
2021PhoneJPG_reduced C+D Contains the jpg images collected with a smartphone in 2021 that have a corresponding white-balanced image. Corresponds to splits C and D.
2021PhoneJPG_WB_reduced C+D Contain all the white-balanced images collected with a smartphone in 2021. Corresponds to splits C and D.
2021PhoneJPG_WB B+C+D Contains all the jpg images collected with a smartphone in 2021, from splits B, C, and D
2021ReflexJPG_WB B+C+D Contains all the jpg images collected with a reflex camera in 2021 from splits B, C, and D
2022PhoneJPG_WB F+H+I Contains all the jpg images collected with a smartphone in 2022 from splits F, H, and I
2022d435iAutoWB F+H+I Contains all the auto white balanced jpg images collected with d435i in 2022, from splits F, H, and I
Table 6: Description of the basic splits used to train and test the proposed algorithms.
Name Train set components Test set components Description
Season21ToSeason22Phone A+C+D F+H+I All jpg images taken with a phone.
Season21ToSeason22Phone_WB B+C+D F+H+I All raw images taken with a phone that have been white-balanced.
Season21ToSeason22All_WB B+C+D F+H+I All phone and reflex images white-balanced.
Phone_d435i_Std A+C+D+F+H+I All phone images not white-balanced, plus all d435i images
AllWB_d435i_Std B+C+D+F+H+I All phone and reflex images white-balanced, plus all d435i images
MixedSeason_JPG A+C+D+H+I F All jpg images, keeping for testing only the ones with the most uncertain brix value (around 16).
MixedSeason_WB_WB B+C+D+H+I F All white balanced images, keeping for test only the most uncertain brix value (around 16).
Table 7: This table describes the most complex data splits, which we called mixed splits. These splits span across different sensors, while the train and test separation is based on the season.
Split name #$\#samples$ $Brix_{mean}$ $Brix_{std}$ $Brix_{mae}$ $CC_{1}$ $CC_{2}$ $CC_{3}$ $CC_{4}$ $CC_{5}$ $CC_{6}$
2021PhoneJPG 212 16.23 2.52 2.02 0.01 0.04 0.09 0.14 0.27 0.45
2021PhoneJPG_WB 150 17.54 2.46 2.04 0.01 0.03 0.06 0.10 0.21 0.59
2021PhoneJPG_reduced 112 17.19 2.43 1.99 0.01 0.04 0.07 0.12 0.22 0.54
2021PhoneJPG_WB_reduced 112 17.19 2.43 1.99 0.01 0.04 0.07 0.12 0.22 0.54
2022PhoneJPG 92 18.91 2.16 1.71 0.01 0.03 0.15 0.16 0.37 0.27
2022PhoneJPG_WB 92 18.91 2.16 1.71 0.01 0.03 0.15 0.16 0.37 0.27
2021ReflexJPG_WB 150 17.54 2.46 2.04 0.01 0.03 0.06 0.10 0.21 0.59
2022d435iAutoWB 92 18.91 2.16 1.71 0.01 0.03 0.15 0.16 0.37 0.27
MixedSeason_train 352 17.67 2.75 2.33 0.01 0.04 0.12 0.16 0.32 0.34
MixedSeason_test 44 15.96 1.68 1.36 0.00 0.00 0.09 0.09 0.27 0.55
MixedSeason_WB_train 440 18.27 2.40 1.98 0.01 0.04 0.10 0.13 0.27 0.46
MixedSeason_WB_test 44 15.96 1.68 1.36 0.00 0.00 0.09 0.09 0.27 0.55
Season21ToSeason22Phone_train 212 16.23 2.52 2.02 0.01 0.04 0.09 0.14 0.27 0.45
Season21ToSeason22Phone_test 92 18.91 2.16 1.71 0.01 0.03 0.15 0.16 0.37 0.27
Season21ToSeason22Phone_WB_train 150 17.54 2.46 2.04 0.01 0.03 0.06 0.10 0.21 0.59
Season21ToSeason22Phone_WB_test 92 18.91 2.16 1.71 0.01 0.03 0.15 0.16 0.37 0.27
Season21ToSeason22All_WB_train 300 17.54 2.46 2.04 0.01 0.03 0.06 0.10 0.21 0.59
Season21ToSeason22All_WB_test 92 18.91 2.16 1.71 0.01 0.03 0.15 0.16 0.37 0.27
All_Std 396 17.48 2.71 2.30 0.01 0.04 0.12 0.15 0.32 0.37
All_Std_WB 484 18.06 2.44 2.03 0.01 0.03 0.10 0.12 0.27 0.47
Table 8: Basic statistics for all the atomic and mixed splits described in Table 7. For each split the following statistics are listed: number of samples, average °Brix, standard deviation of the °Brix, °Brix mean absolute error, and the percentage of each color class in the dataset. The color classes are unbalanced towards the unripe color classes, since the ripe bunches were harvested every few days.
Appendix B Hyper-parameters Exploration for the Histogram based Estimators
Following from Section 3.2, here we give the details of the hyper-parameter search for the histogram features.
Table 9 summarizes the hyper-parameter ranges investigated by this experiment. Most of them have been introduced in Section 2.2. In addition to those, here we add the investigation of the effect of the input image resolution (low_res: $320\times 240$, med_res: $640\times 480$, high_res: $1280\times 720$). For the $cross$ parameter, to avoid hyper-parameter explosion, we designed three possible configurations (none, thin, fat), as described in Section 2.2.
@@ Appendix B Hyper-parameters Exploration for the Histogram based Estimators
$\alpha$ 0.33, 1, 3, 9, 27, 81, 243, 729 Ridge Regression regularization parameter.
Table 9: Hyper-parameters ranges for the histogram feature based models.
To identify the most relevant hyper-parameters, taking into account the non-linear interactions between the different features, we opted to train a random forest (RF) algorithm on the hyper-parameter values themselves, to predict the validation error. Using this technique allows for reading the importance score of the RF as an indicator of the relevance of each hyper-parameter [5]. Figure 11(a) shows the importance factors for the RF regressor trained to predict the validation error from the hyper-parameter values. The importance factor show which features contributes the most to the RF regression of the validation error, allowing for concluding that the different starting image resolutions have little influence on the SSC estimation. This is a relevant finding, since it allow us to safely work with the lowest resolution images without trading for estimation accuracy. Among the most important variables we find the ones that control the model flexibility, i.e. the regularization parameter, the $x$ and $y$ bin number, and the cross parameters, that impact the feature vector length. Together these parameters have to balance prediction power with the risk of over-fitting.
Finally, the color space show some relevance too. The correlation coefficient has already been shown in Fig. 8(a), and its replicated in Fig. 11(b) for convenience. Image resolution was omitted and the total histogram length ($feature\_len$) was added. That histogram helps understanding which variable linearly correlates with the validation error, in positive or negative way, but alone it’s not enough to show feature importance, since there could be some interactions effects between them.
@@
Figure 11: a) Training a Random Forest using the hyper-parameters values to predict the validation error and then extracting the importance scores is a common way to extract the relative impact of each feature on the final estimator performance. Here we can see that the choice of initial image resolution is not relevant, while the regularization, the number of bins (that influence feature length and flexibility), and the color space have to carefully selected. b) shows the correlation computed on the 10% best performing runs. This plot complements the other, showing where the correlation with the error is positive or negative.
Figure 12: To better extract the impact of feature length on validation error, we add three parameters to the analysis, $feature_{l}en$ and its interactions with $alpha$ and $\log{alpha}$. It is easy to see that the length of the histogram is the most important hyper-parameter to be tuned for final model performance.
To further explore this influence of feature flexibility due to histogram length, we show in Figure 12 the correlation between the validation error and the most important hyper-parameters extracted from Figure 11(a), explicitly adding $feature\_lenght$ and its interactions with $alpha$ and $\log(alpha)$. It is evident that the feature length is the most important factor among the others, and that the interaction between feature length and alpha is more than relevant. The general behavior that we can extract from this analysis is that the more are the bins of the histogram, the higher should be the alpha value, and together these two parameters are sufficient to explain a great deal of the validation error variability.
Appendix C Hyper-parameters Exploration for the Multi-task DNN
This appendix follow the same principles described in the previous one. Table 10 gives a summary of the hyper-parameters ranges explored in the sweep. Given the huge parameter space, this time the sweep was performed with random sampling, with a total run count of 4912 (see Section 3.3). Figures 13(a) and 14(a) show importance values for the RF trained on the hyper-parameter values to predict validation BCH respectively for $T2_{STD}$ (Season21ToSeason22_Phone) and $T3_{STD}$ (MixedSeason). Finally, Tables 4, 11, 13, 14, and 12 summarize the best 10 runs respectively on $T2_{STD}$, $T2_{WB}$, $T3_{STD}$, $T3_{WB}$ and $T2_{ALL}$.
Multi-task learning hyper-parameter Range description
$\alpha$ and $\beta$ 1, 10, 50, 100, 500, 1000 The loss balancing factors (Eq. 2)
$batch\_size$ 32, 64, 128 Training batch size
$learning\_rete$ 0.00001, 0.00005, 0.000001 Initial learning rate, to be modified by the scheduling strategy
$learning\_rete$ 0.00001, 0.00005, 0.000001 Initial learning rate, to be modified by the scheduling strategy
$momentum$ 0.8, 0.9, 0.95, 0.99 Momentum for the SGD optimizer
@@ Appendix C Hyper-parameters Exploration for the Multi-task DNN
$one\_cycle$ $True$ or $False$ Whether to use of not the one cycle scheduling strategy
$pretrained$ $True$ or $False$ Whether the resnet18 backbone is pretrained on ImageNet or not.
$dataset\_split$ $Season21ToSeason22Phone$, $Season21ToSeason22Phone\_WB$, $Season21ToSeason22All\_WB$, $MixedSeason$, $MixedSeason\_WB$ Image resolution, corresponding to $320\times 240$, $640\times 480$ and $1280\times 720$ respectively, in color corrected and standard form.
$\alpha$ 0.33, 1, 3, 9, 27, 81, 243, 729 Ridge Regression regularization parameter.
@@
Table 10: This Table summarizes the hyper-parameters descriptions for the Multi-task experiments. A wide range of hyper-parameters was explored using random search for a total of more than $5000$ experimental runs. In this case, the dataset splits used are not to be considered a proper hyper-parameter, but as a defining subset of the parameter sweep, since the distribution shifts between them are considerable. They are included in this table for ease of reference.
Figure 13: a) These plots show the importance and correlation factor of the best $25\%$ of the runs trained on the $Season2021ToSeason2022$ dataset split. It can be noted that the main influence comes from the classical parameters that impact any DNN training (learning rate, batch size, regularization parameters), while the choice of image size and color space comes seems to be still relevant for the final performances. More insights on this can be extracted from Table 4.
Figure 14: These plots show the importance and correlation factor of the best $25\%$ of the runs trained on the $MixedSeason$ dataset split.
Val BCH Val SSC MAE Val Acc $alpha$ $beta$ $im\_width$ $im\_height$ $momentum$ $learning\_rate$ $batch\_size$ $blur\_\sigma\_min$ $blur\_\sigma\_max$ $pretrained$ $one\_cycle$ $HSV$ $Lab$ $augment\_extra$
4.100699 1.835277 0.478261 500 50 128 256 0.800000 0.000050 64 0.500000 2 True True True False True
4.132786 1.913489 0.576087 100 1 256 64 0.990000 0.000050 64 0.010000 2 True True False False True
4.384027 1.570156 0.554348 1000 10 256 256 0.990000 0.000001 32 0.010000 2 True False False False True
4.397698 1.605315 0.434783 500 10 128 128 0.800000 0.000050 32 0.500000 2 False True False True True
4.499346 1.982949 0.576087 100 100 256 256 0.950000 0.000050 64 0.010000 4 True True True False False
4.522784 1.714363 0.500000 1000 100 256 128 0.990000 0.000010 64 0.100000 2 True False False True False
4.552584 1.768912 0.489130 500 10 128 128 0.900000 0.000050 64 0.500000 2 True True False True False
4.558001 1.644159 0.445652 1000 500 256 128 0.800000 0.000010 32 0.100000 4 False True False True False
4.583944 1.916520 0.521739 500 10 256 256 0.950000 0.000010 64 0.500000 4 False False False False True
4.596296 1.968404 0.597826 100 1 64 64 0.990000 0.000050 32 0.100000 4 True True True False True
Table 11: The best ten runs on $Season2021ToSeason2022\_WB$ split. Compared to the ones given in Table 4, both the SSC MAE estimation error and Color estimation accuracy show worse performance.
Val BCH Val SSC MAE Val Acc $alpha$ $beta$ $im\_width$ $im\_height$ $momentum$ $learning\_rate$ $batch\_size$ $blur\_\sigma\_min$ $blur\_\sigma\_max$ $pretrained$ $one\_cycle$ $HSV$ $Lab$ $augment\_extra$
3.977163 1.754222 0.576087 50 1 128 128 0.990000 0.000050 32 0.500000 2 True True False True True
4.123796 1.945782 0.532609 100 10 128 128 0.950000 0.000050 64 0.010000 2 False False True False False
4.269722 1.610763 0.673913 500 500 256 256 0.990000 0.000010 64 0.100000 4 True True True False True
4.307943 1.754948 0.608696 500 10 128 256 0.950000 0.000010 32 0.100000 2 True True False False False
4.327728 1.697267 0.467391 1000 50 128 128 0.950000 0.000010 32 0.500000 4 False True False True False
4.336076 1.930476 0.532609 1000 10 64 256 0.800000 0.000050 64 0.010000 2 True False False True False
4.347323 1.688515 0.500000 500 100 256 256 0.900000 0.000010 64 0.100000 4 False False False True True
4.411548 1.531855 0.597826 500 50 256 256 0.990000 0.000010 32 0.500000 2 False False False False False
4.440626 1.685545 0.630435 1000 500 256 256 0.990000 0.000001 128 0.100000 2 False False True False False
4.501309 1.803873 0.608696 1000 10 256 256 0.800000 0.000050 64 0.100000 4 True False False True True
Table 12: The best ten runs on $Season2021ToSeason2022All\_WB$ split.
Val BCH Val SSC MAE Val Acc $alpha$ $beta$ $im\_width$ $im\_height$ $momentum$ $learning\_rate$ $batch\_size$ $blur\_\sigma\_min$ $blur\_\sigma\_max$ $pretrained$ $one\_cycle$ $HSV$ $Lab$ $augment\_extra$
1.581554 1.200569 0.795455 500 10 128 64 0.900000 0.000050 32 0.010000 2 True False True False True
1.662561 1.116875 0.772727 1000 50 256 128 0.990000 0.000010 32 0.100000 4 True False False False True
1.686950 1.162343 0.772727 1000 50 256 256 0.950000 0.000010 32 0.010000 4 True True False True False
1.723328 1.231537 0.795455 50 100 256 256 0.950000 0.000010 32 0.010000 4 False False True False True
1.741897 1.102953 0.795455 1000 1000 64 64 0.990000 0.000001 32 0.500000 4 False True False True True
1.745264 1.301862 0.795455 1000 1000 128 128 0.950000 0.000001 32 0.100000 4 False True False False False
1.746910 1.269346 0.818182 1000 1000 256 256 0.900000 0.000010 128 0.100000 2 True True False True False
1.771968 1.174211 0.795455 100 500 64 128 0.950000 0.000010 64 0.100000 4 False False False False True
1.788137 1.190380 0.772727 100 50 128 256 0.950000 0.000050 128 0.500000 4 False False False False False
1.788171 1.242672 0.818182 10 10 256 256 0.950000 0.000050 64 0.100000 4 True True False False True
Table 13: The best ten runs on $MixedSeason$ ($T3_{STD}$) split. The validation set is the same as for the $MixedSeason\_WB$ ($T3_{WB}$), that is collected with a d435i camera. Since the images from this device are automatically white balanced, this imply that the model faces a strong covariate shift due to different camera device and image modality.
Val BCH Val SSC MAE Val Acc $alpha$ $beta$ $im\_width$ $im\_height$ $momentum$ $learning\_rate$ $batch\_size$ $blur\_\sigma\_min$ $blur\_\sigma\_max$ $pretrained$ $one\_cycle$ $HSV$ $Lab$ $augment\_extra$
1.545861 1.307749 0.818182 50 100 64 128 0.800000 0.000050 32 0.500000 4 True True False False False
1.596429 1.101236 0.795455 100 100 128 256 0.800000 0.000010 128 0.010000 2 False False True False False
1.611784 1.219757 0.772727 500 1000 128 128 0.800000 0.000010 32 0.010000 4 False False True False False
1.639119 1.065401 0.727273 50 10 128 256 0.950000 0.000050 64 0.500000 2 False True False False True
1.688067 1.220433 0.750000 500 500 128 128 0.950000 0.000010 32 0.500000 4 True False True False False
1.709684 1.295742 0.704545 10 10 256 256 0.990000 0.000010 32 0.010000 2 True False False True False
1.715017 1.159092 0.772727 100 100 64 128 0.800000 0.000050 32 0.100000 2 True False True False False
1.715871 1.158179 0.840909 10 10 256 256 0.900000 0.000050 64 0.100000 2 True False True False True
1.759702 1.200407 0.795455 50 50 128 128 0.900000 0.000050 128 0.100000 4 False False True False False
1.776299 1.200725 0.772727 100 50 256 256 0.900000 0.000050 32 0.010000 4 True True True False False
Table 14: The best ten runs on $MixedSeason\_WB$ ($T3_{WB}$)split. This time the results are comparable with the STD ones presented in Table 13. This can be explained considering the fact that the validation images of the d435i are white-balanced and therefore with a similar hue to the color-corrected training images.

### v2

@@ Acknowledgments
This work is part of a project that has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 101016906 – Project CANOPIES
@@
Appendix A Hyper-parameters Exploration for the
Histogram-based Estimators
Following from Section 3.2, to allow full reproducibility of the experiments, here we give the details of the hyper-parameter search for the histogram features.
Table 8 summarizes the hyper-parameter ranges investigated in this experiment.
Table 8: Hyper-parameters ranges for the histogram feature-based models.
@@ Appendix A Hyper-parameters Exploration for the
$\lambda$ 0.33, 1, 3, 9, 27, 81, 243, 729 Ridge Regression regularization parameter.
Given the manageable number of combinations and the low computational cost of each experiments, we performed a full grid search to find the optimal hyper-parameters. To identify the most relevant ones we used the importance factor, as described in Grömping, [5] and implemented in the Weights&Biases software suite for experiments tracking [3]. Fig. 11 gives an example of this analysis, while Table 9 shows some of the best performing runs. The full results and hyper-parameters for all 3888 experimental runs are publicly available for inspection to ensure full transparency in the GitHub11footnotemark: 1 and Weights&Biases [^5] https://wandb.ai/sparviero19/CQEM-193_Brix_sweep/reports/CQEM-193-Sweep-results-on-sweep-zgyffmc5--Vmlldzo2Njc3MjA4 companion sites.
Table 9: Top 10 histogram-based models, ranked by validation mean absolute error (Val MAE), along with their corresponding hyper-parameters values.
Test MAE Val MAE Train MAE $\lambda$ $n_{bin_{x}}$ $n_{bin_{y}}$ $cross$ $color\_space$
1.46 2.10 1.58 9 16 8 none HSV
1.51 2.13 1.55 9 8 16 none HSV
1.52 2.14 1.61 3 8 8 none HSV
1.51 2.14 1.18 9 8 32 none HSV
1.52 2.16 1.62 1 16 8 thin CIE-Lab
1.49 2.16 1.25 3 16 8 none HSV
1.52 2.18 1.43 3 16 8 fat HSV
1.52 2.19 1.72 0.33 8 8 thin CIE-Lab
1.43 2.20 1.19 3 8 16 none HSV
1.49 2.33 1.22 1 16 8 thin HSV
@@
Figure 11: Training a Random Forest using the hyper-parameter values to predict the validation error and then extracting the importance scores is a common way to extract the relative impact of each feature on the final estimator performance. Here we can see that the choice of initial image resolution is not relevant, while the regularization, the number of bins (that influence feature length and flexibility), and the color space have to be carefully selected.
Table 10: Hyper-parameter descriptions for the Multi-task experiments.
Hyper-parameter name Range description
$\alpha$ and $\beta$ 1, 10, 50, 100, 500, 1000 The loss balancing factors (Eq. 2).
$batch\_size$ 32, 64, 128 Training batch size.
$learning\_rete$ 0.00001, 0.00005, 0.000001 Initial learning rate, to be modified by the scheduling strategy.
$momentum$ 0.8, 0.9, 0.95, 0.99 Momentum for the SGD optimizer.
@@
$one\_cycle$ $True$ or $False$ Whether to use of not the one cycle scheduling strategy (Smith and Topin, [16]).
$pretrained$ $True$ or $False$ Whether the resnet18 backbone is pretrained on ImageNet (He et al., [7]) or not.
[…]
Appendix B Hyper-parameter Exploration for the Multi-task DNN
@@
This appendix follows the same principles described in the previous one. Table 10 gives a summary of the hyper-parameter ranges explored in the sweep. A wide range of hyper-parameters was explored using random search. In this case, the dataset splits used are not to be considered proper hyper-parameters, but as a defining subset of the parameter sweep, since the distribution shifts between them are considerable. They are included in this table for ease of reference.
Given the huge parameter space, this time the sweep was performed with random sampling, with a total run count greater than 4000. To understand which parameters were most influential, we analyzed the top 25% of all experimental runs. An example of the correlation between some of the hyper-parameters and the BCH validation score for this subset is visualized in Fig. 12. The full analysis revealed that standard deep learning parameters like learning rate and regularization weights, along with image dimensions, were the most influential factors on performance. The detailed hyper-parameter configurations for the best models are available in our GitHub repository and the complete list of experimental run can be found on the Weights&Biases report.222 https://api.wandb.ai/links/sparviero19/s2p10p3a
[…]
Appendix C Multi-Task DNN Hyper-parameter Validation Metric
Our evaluation criteria also align with agronomic quality standards, emphasizing the importance of avoiding the harvest of unripe bunches, while allowing for some variation in quality among ripe bunches. This is achieved by introducing hinge-like penalties that apply only when estimations fall below predefined quality thresholds for SSC ($t_{brix}$) and color ($t_{color}$) estimation, ensuring that models are primarily evaluated on their ability to meet minimum quality standards. The final score, $\mathcal{S}$ integrates the normalized performance metrics and the agronomic quality criteria. We call this metric Brix-Color-Hinge (BCH) metric.
The formal definition of the score $\mathcal{S}$ is the following:
$L_{BH}$ $=\max(0,L_{brix}-t_{brix}))^{\rho}$ (4)
$L_{CH}$ $=\max(0,-\eta(A_{color}-t_{color}))^{\rho}$ (5)
where $\rho$ is the degree of the two hinge penalties $\mathcal{L}_{BH}$ (for °Brix) and $\mathcal{L}_{CH}$ (for color), set to 2 in all our experiments. $t_{brix}$ and $t_{color}$ represent the agronomic quality thresholds for °Brix and color, respectively, and $\eta$ is the normalization factor between SSC and color scores. In our experiments, $t_{brix}=2.5$Brix, which is comparable with the standard deviation of the training data, and $t_{color}=0.8$, since an error rate of $80\%$ is usually considered acceptable for human performance in this context The Canopies Consortium, [17]. The normalization factor is empirically set as
$\eta=\frac{t_{brix}}{t_{color}}.$ (6)